OpenStudy (anonymous):

taking exam please help. will give medal x^2 =50 plus or minus square root 50 plus or minus square root 100 ±25 ±100

pooja195 (pooja195):

We dont help on exam questions thats cheating but i'll tell you this much \[\huge~\rm~\sqrt{x^2}=\sqrt{50}\]

pooja195 (pooja195):

Thats all im saying.

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

its an online test. please help me @pooja195

pooja195 (pooja195):

and? It doesnt mean u can cheat. I gave you a clue.

pooja195 (pooja195):

Whats the square root of 50 ? whats the square root of x^ thats all it is.

OpenStudy (anonymous):

well I'm thinking its the 3rd one am I right @pooja195

pooja195 (pooja195):

yes! :D

OpenStudy (anonymous):

thank you

pooja195 (pooja195):

:)